# Copula Conformal Prediction for Multi-step Time Series Forecasting

### 23 Jan 2024 v3

@@ B.1 Upper and Lower bounds for Copulas
To provide a better understanding of the properties of Copulas, consider the Frechet-Hoeffding Bounds (Theorem B.1). In fact, the Frechet-Hoeffding upper- and lower- bounds are both copulas. The lower bound is percisely the Bonferroni correction used in Stankevičiūtė et al. (2021) - therefore by estimating the copula more precisely instead of using a lower bound, we have a guaranteed efficiency improvement for the confidence region.
@@ B.2 Numerical optimization with SGD for search
The indicator function is implemented using a sigmoid function whose input is multiplied by a constant for differentiablility. A small amount of L2 regularization is added to the loss function to ensure the searched scores are as low as possible. We use the Adam optimizer and perform gradient descent for 500 steps to get the final result. The optimization process to find $\mathbf{s}^{*}$ typically takes a few seconds on CPU. For each run of our experiments, the calibration and prediction steps of CopulaCPTS combined took less than 1 minute to run on an Apple M1 CPU. Please refer to the CP class in reference code for implementation details.
@@ B.3 CopulaCPTS in auto-regressive forecasting
Auto-regressive forecasting is a common framework in time series forecasting. So far, we have been looking at forecasts for a predetermined number of time steps $k$. One can use a fixed length model to forecast variable horizons $k^{\prime}$ autoregressively, taking the model output as part of the input. In the conformal prediction setting, we want not only to autoregressively use the point forecasts, but also to propagate the uncertainty measurement.
If the time series and uncertainty are stationary (for example: additive Gaussian noise), the copula remains the same for any sliding window of $k$ steps, i.e. $C(u_{1},\ldots,u_{k})=C(u_{2},\ldots,u_{k+1})$. Therefore, after finding $(u_{1}^{*},\ldots,u_{k}^{*})$ such that $C(u_{1}^{*},\ldots,u_{k}^{*})\geq 1-\alpha$, we can simply search for $u_{k+1}^{*}$ such that $C(u_{2}^{*},\ldots,u_{k}^{*},u_{k+1}^{*})\geq 1-\alpha$. The guarantee proven in Theorem 4.1 still holds for the new estimate. In this way, we can achieve the coverage guarantee over the entire autoregressive forecasting horizon.
@@ B.4 Autoregressive prediction
We now provide a toy scenario to illustrate when re-estimating the copula is necessary and improves validity. Consider a time series of three time steps $t_{0},t_{1},t_{2}$. The two scenarios are illustrated in Figure 6. In both scenarios the mean and variance of all timeseps are 00 and $1$ respectively. In scenario (a), $t_{0}=t_{1}$ and hence their covariance is $1$. The copula estimated on $t_{0}$ and $t_{1}$ is $C_{0:1}(F(t_{0}),F(t_{1}))=F(t_{0})=F(t_{1})$. This copula will significantly underestimate the confidence region of $t_{2}$ where its covariance with $t_{1}$ is $-1$. In fact the coverage of $C_{0:1}(F_{1}(t_{1}),F_{2}(t_{2}))=0.74$. On the other hand, (b) illustrates a scenario where the copula for any 2 consecutive time series remain the same $C_{0}=C_{1}$. In this case, applying $C_{0}$ directly to forecast $C_{1}$ achieves precisely 90% coverage.
@@ C.1 Underlying forecasting models
Particle Dataset
The underlying forecasting model for the particle experiments is an 1-layer LSTM network with embedding size $=24$. The hidden state is then passed through a linear network to forecast the timesteps concurrently (output has dimension $k\times d_{y}$). We train the model for 150 epochs with batch size 128. Hyperparameters of the network are selected through a model search by performance on a 5-fold cross validation split of the dataset. The architecture and hyperparameters are shared for all baselines and CopulaCPTS in Table 1.
Drone
For the drone trajectory forecasting task, we use he same LSTM forecasting network as the particle dataset, but with hidden size increased to 128. We train the model for 500 epochs with batch size 128. The same architecture and hyperparameters are shared for all baselines and CopulaCPTS reported in Table 1.
Covid-19
The base forecasting model for Covid-19 dataset is the same as the synthetic datasets, with hidden size $=128$ and were trained for 150 epochs with batch size 128. The same architecture and hyperparameters are shared for all baselines and CopulaCPTS reported in Table 1.
Argoverse
As highlighted in the main text, we utilize a state-of-the-art prediction algorithm LaneGCN (Liang et al., 2020) as the underlying forcaster model for CF-RNN and Copula-RNN. We refer the readers to their paper and code base for model details. The architecture of the RNN network used for MC-Dropout and BJRNN is an Encoder-Decoder network. Both the encode and decoder contains a LSTM layer with encoding size 8 and hidden size 16. We chose this architecture because the is part of the official Argoverse baselines (https://github.com/jagjeet-singh/argoverse-forecasting) and demonstrates competitive performance.
C.2 COVID-19 Dataset
The COVID-19 dataset is downloaded directly from the official UK government website https://coronavirus.data.gov.uk/details/download by selecting region for area type and newCasesByPublishDate for metric. There are in total 380 regions and over 500 days of data, depending when it is downloaded. We selected 150-day time series from the collection to construct our dataset.
C.3 Calibration and Efficiency chart for COVID-19
Figure 7 shows comparison of calibration and efficiency for the daily new COVID 19 cases forecasting.
Figure 7: Calibration and efficiency comparison on different $\epsilon$ level for COVID-19 Daily Forecasts. The copula methods (orange and red lines) are more calibrated (coinciding with the green doted line) and sharp (low width) compared to baselines.
To see if the daily fluctuation due to testing behaviour disrupts other method, we also ran the same experiment on weekly aggregated new cases forecast. We take 14 weeks of data as input and output forecasts for the next 6 weeks. The results are illustrated in Figure 8. The weekly forecasting scenario gives us similar insights as the daily forecasts.
@@ C.4 Argoverse
C.5 Additional Experiment Results
We present in figure 8 and 9 some qualitative results for uncertainty estimation.
To test how the effects of copulaCPTS compare with baseline on other base forecasters, we also include an encoder-decoder architecture with the same embedding size as the RNN models introduced in appendix C.1 for each dataset. The results are presented in Table 3. We omit these results in the main text because we found that they do did not bring significant improvement to time series forecasting UQ.
@@ C.5 Additional Experiment Results
Figure 9: $99\%$ Confidence region produced by three methods for the drone dataset. Copula methods (a) produces a more consistent, expanding cone of uncertainty compared to MC-Dropout (b) sharper one compared to CF-RNN (c).
Figure 10: Illustrations for confidence regions given by CF-RNN (blue) and CopulaCPTS (orange) in at time steps 0, 10, 20, and 30. Note that in order to achieve 90% coverage, the regions are larger than needed, especially in straight-lane cases like the middle two. Using copulas to couple together time steps results in a much smaller region, while achieving similarly good coverage.
@@
Table 3: Additional results. Copula methods achieves high level of calibration while producing sharper predictions regions. The sharpness gain is even more pronounced in at higher confidence levels (99%), where we want the prediction region to be useful while remaining valid.
@@ 1 Step 5 Steps 15 Steps
C.6 Study on $\alpha_{j}$ search
@@
The $\alpha_{j}$ values are an indicator of how interrelated the uncertainty between each time step are: Bonferroni Correction used in Stankevičiūtė et al. (2021) (grey dotted line in Figure 11) assumes that the time steps are independent, with CopulaCPTS we have lower $1-\alpha_{j}$ levels while having valid coverage (blue and orange lines in Figure 11). This shows that the uncertainty of the time steps are not independent, and we are able to utilize this dependency to shrink the confidence region and still maintaining the coverage guarantee.
Table 5 shows that there are no significant difference between coverage and area performance for the two search methods with in the scope of datasets we study in this paper. However, we want to highlight that SGD search is $O(n)$ complexity to optimization steps, regardless to the prediction horizon. SGD also allows for varying $\alpha_{j}$ which might be useful in some settings, for example capturing uncertainty spikes for some time steps as seen in the COVID-19 dataset of Figure 11. Dichotomy search, on the other hand, is $O(nlog(n))$ complexity to the search space depends on granularity, and will be $O(knlog(kn)$ if we want to search for varying $\alpha_{j}$.
@@
Table 5: Coverage and area comparison between stochastic search for fixed $\alpha_{j}$ and SGD for Varying $\alpha_{j}$. We do not see significant difference between the performance of two.
@@
C.7 Comparison to additional Baselines
@@
L2-Conformal. L2-Conformal uses the same underlying RNN forecaster as CF-RNN and Copula RNN. We use that nonconformity score of the vector norm of all timesteps concatenated together $\|\hat{\textbf{y}}_{t+1:t+k}-\textbf{y}_{t+1:t+k}\|$ to perform ICP. As there are no analytic way to represent a $k\times d_{y}$-dimensional uncertainty region on 2-D space, we calculate the area and plot the region for L2 Conformal baseline with the maximum deviation at each timestep such that the vector norm still stays within range.
Direct Gaussian. Direct Gaussian use the same model architecture and training hyperparameters, with the addition of a linear layer that outputs the variance for each timestep, and is optimized using negative log loss, a proper scoring rule for probabilistic forecasting. We obtain the area by analytically calculating the 90% confidence interval for each variable.
Results in Table 6 shows that L2-conformal produces inefficient confidence area, and directly outputting variance under-covers test data. These results align with previous findings and motivates our method, which is both more calibrated and sharper compared to these baselines. We show a visualization in Figure 12 to illustrate the different properties of the methods qualitatively.

### 18 Mar 2024 v4

@@ B.1 Upper and Lower bounds for Copulas
To provide a better understanding of the properties of Copulas, consider the Frechet-Hoeffding Bounds (Theorem B.1). In fact, the Frechet-Hoeffding upper- and lower- bounds are both copulas. The lower bound is precisely the Bonferroni correction used in Stankevičiūtė et al. (2021) - therefore by estimating the copula more precisely instead of using a lower bound, we have a guaranteed efficiency improvement for the confidence region.
@@ B.2 Numerical optimization with SGD for search
The indicator function is implemented using a sigmoid function whose input is multiplied by a constant for differentiability. A small amount of L2 regularization is added to the loss function to ensure the searched scores are as low as possible. We use the Adam optimizer and perform gradient descent for 500 steps to get the final result. The optimization process to find $\mathbf{s}^{*}$ typically takes a few seconds on CPU. For each run of our experiments, the calibration and prediction steps of CopulaCPTS combined took less than 1 minute to run on an Apple M1 CPU. Please refer to the CP class in the reference code for implementation details.
@@ B.3 CopulaCPTS in auto-regressive forecasting
Auto-regressive forecasting is a common framework in time series forecasting. So far, we have been looking at forecasts for a predetermined number of time steps $k$. One can use a fixed-length model to forecast variable horizons $k^{\prime}$ autoregressively, taking the model output as part of the input. In the conformal prediction setting, we want not only to autoregressively use the point forecasts, but also to propagate the uncertainty measurement.
If the time series and uncertainty are stationary (for example additive Gaussian noise), the copula remains the same for any sliding window of $k$ steps, i.e. $C(u_{1},\ldots,u_{k})=C(u_{2},\ldots,u_{k+1})$. Therefore, after finding $(u_{1}^{*},\ldots,u_{k}^{*})$ such that $C(u_{1}^{*},\ldots,u_{k}^{*})\geq 1-\alpha$, we can simply search for $u_{k+1}^{*}$ such that $C(u_{2}^{*},\ldots,u_{k}^{*},u_{k+1}^{*})\geq 1-\alpha$. The guarantee proven in Theorem 4.1 still holds for the new estimate. In this way, we can achieve the coverage guarantee over the entire autoregressive forecasting horizon.
@@ B.4 Autoregressive prediction
We now provide a toy scenario to illustrate when re-estimating the copula is necessary and improves validity. Consider a time series of three time steps $t_{0},t_{1},t_{2}$. The two scenarios are illustrated in Figure 6. In both scenarios, the mean and variance of all time steps are 00 and $1$ respectively. In scenario (a), $t_{0}=t_{1}$ and hence their covariance is $1$. The copula estimated on $t_{0}$ and $t_{1}$ is $C_{0:1}(F(t_{0}),F(t_{1}))=F(t_{0})=F(t_{1})$. This copula will significantly underestimate the confidence region of $t_{2}$ where its covariance with $t_{1}$ is $-1$. In fact the coverage of $C_{0:1}(F_{1}(t_{1}),F_{2}(t_{2}))=0.74$. On the other hand, (b) illustrates a scenario where the copula for any 2 consecutive time series remains the same $C_{0}=C_{1}$. In this case, applying $C_{0}$ directly to forecast $C_{1}$ achieves precisely 90% coverage.
@@ C.1 Underlying forecasting models
Particle Dataset.
The underlying forecasting model for the particle experiments is an 1-layer LSTM network with embedding size $=24$. The hidden state is then passed through a linear network to forecast the timesteps concurrently (output has dimension $k\times d_{y}$). We train the model for 150 epochs with batch size 128. Hyperparameters of the network are selected through a model search by performance on a 5-fold cross-validation split of the dataset. The architecture and hyperparameters are shared for all baselines and CopulaCPTS in Table 1.
Drone.
For the drone trajectory forecasting task, we use the same LSTM forecasting network as the particle dataset, but with its hidden size increased to 128. We train the model for 500 epochs with batch size 128. The same architecture and hyperparameters are shared for all baselines and CopulaCPTS reported in Table 1.
Covid-19.
The COVID-19 dataset is downloaded directly from the official UK government website https://coronavirus.data.gov.uk/details/download by selecting region for area type and newCasesByPublishDate for metric. There are in total 380 regions and over 500 days of data, depending on when it is downloaded. We selected 150-day time series from the collection to construct our dataset.
The base forecasting model for Covid-19 dataset is the same as the model for synthetic datasets, with hidden size $=128$, and were trained for 150 epochs with batch size 128. The same architecture and hyperparameters are shared for all baselines and CopulaCPTS reported in Table 1.
Argoverse.
As highlighted in the main text, we utilize a state-of-the-art prediction algorithm LaneGCN (Liang et al., 2020) as the underlying forecaster model for CF-RNN and Copula-RNN. We refer the readers to their paper and code base for model details. The architecture of the RNN network used for MC-Dropout and BJRNN is an Encoder-Decoder network. Both the encode and decoder contain a LSTM layer with encoding size 8 and hidden size 16. We chose this architecture because the is part of the official Argoverse baselines (https://github.com/jagjeet-singh/argoverse-forecasting) and demonstrates competitive performance.
C.2 Calibration and Efficiency chart for COVID-19
Figure 7 shows a comparison of calibration and efficiency for the daily new COVID 19 cases forecasting.
Figure 7: Calibration and efficiency comparison on different $\epsilon$ level for COVID-19 Daily Forecasts. The copula methods (orange and red lines) are more calibrated (coinciding with the green dotted line) and sharp (low width) compared to baselines.
To see if the daily fluctuation due to testing behavior disrupts other methods, we also ran the same experiment on weekly aggregated new cases forecast. We take 14 weeks of data as input and output forecasts for the next 6 weeks. The results are illustrated in Figure 8. The weekly forecasting scenario gives us similar insights as the daily forecasts.
@@ C.3 Argoverse
C.4 Additional Experiment Results
We present in Figures 8 and 9 some qualitative results for uncertainty estimation.
To test how the effects of copulaCPTS compare with baseline on other base forecasters, we also include an encoder-decoder architecture with the same embedding size as the RNN models introduced in Appendix C.1 for each dataset. The results are presented in Table 3. We omit these results in the main text because we found that they did not bring significant improvement to time series forecasting UQ.
@@ C.4 Additional Experiment Results
Figure 9: $99\%$ Confidence region produced by three methods for the drone dataset. Copula methods (a) produce a more consistent, expanding cone of uncertainty compared to MC-Dropout (b) sharper one compared to CF-RNN (c).
Figure 10: Illustrations for confidence regions given by CF-RNN (blue) and CopulaCPTS (orange) at time steps 0, 10, 20, and 30. Note that in order to achieve 90% coverage, the regions are larger than needed, especially in straight-lane cases like the middle two. Using copulas to couple together time steps results in a much smaller region while achieving similarly good coverage.
@@
Table 3: Additional results. Copula methods achieve a high level of calibration while producing sharper prediction regions. The sharpness gain is even more pronounced at higher confidence levels (99%), where we want the prediction region to be useful while remaining valid.
@@ 1 Step 5 Steps 15 Steps
C.5 Study on $\alpha_{j}$ search
@@
The $\alpha_{j}$ values are an indicator of how interrelated the uncertainty between each time step are: Bonferroni Correction used in Stankevičiūtė et al. (2021) (grey dotted line in Figure 11) assumes that the time steps are independent, with CopulaCPTS we have lower $1-\alpha_{j}$ levels while having valid coverage (blue and orange lines in Figure 11). This shows that the uncertainty of the time steps is not independent, and we are able to utilize this dependency to shrink the confidence region while still maintaining the coverage guarantee.
Table 5 shows that there are no significant differences between coverage and area performance for the two search methods within the scope of datasets we study in this paper. However, we want to highlight that SGD search is $O(n)$ complexity to optimization steps, regardless of the prediction horizon. SGD also allows for varying $\alpha_{j}$ which might be useful in some settings, for example capturing uncertainty spikes for some time steps as seen in the COVID-19 dataset of Figure 11. Dichotomy search, on the other hand, is $O(nlog(n))$ complexity to the search space depends on granularity, and will be $O(knlog(kn)$ if we want to search for varying $\alpha_{j}$.
@@
Table 5: Coverage and area comparison between stochastic search for fixed $\alpha_{j}$ and SGD for Varying $\alpha_{j}$. We do not see a significant difference between the performance of the two.
@@
C.6 Comparison to additional Baselines
@@
L2-Conformal. L2-Conformal uses the same underlying RNN forecaster as CF-RNN and Copula RNN. We use the nonconformity score of the vector norm of all timesteps concatenated together $\|\hat{\textbf{y}}_{t+1:t+k}-\textbf{y}_{t+1:t+k}\|$ to perform ICP. As there are no analytic way to represent a $k\times d_{y}$-dimensional uncertainty region on 2-D space, we calculate the area and plot the region for L2 Conformal baseline with the maximum deviation at each timestep such that the vector norm still stays within range.
Direct Gaussian. Direct Gaussian uses the same model architecture and training hyperparameters, with the addition of a linear layer that outputs the variance for each timestep, and is optimized using negative log loss, a proper scoring rule for probabilistic forecasting. We obtain the area by analytically calculating the 90% confidence interval for each variable.
Results in Table 6 show that L2-conformal produces inefficient confidence area, and directly outputting variance under-covers test data. These results align with previous findings and motivate our method, which is both more calibrated and sharper compared to these baselines. We show a visualization in Figure 12 to illustrate the different properties of the methods qualitatively.